OpenStudy (anonymous):

simplify the square root of 3 * square root of 75

OpenStudy (anonymous):

you mean this \[3\times \sqrt{75}\]

OpenStudy (anonymous):

no the square root of 3 times square root of 75

OpenStudy (anonymous):

75=3*5^2 \[\sqrt{3} \times \sqrt{75} = \sqrt{3^{2} \times 5^{2}}=3 \times 5\]

OpenStudy (anonymous):

thanks